OpenStudy (anonymous):

A farmer wants to fence a rectangular garden using 120ft of fencing. The side of the garage will be used for 1 side of the rectangle. If the garden is x feet wide, express the area as a function of x.

OpenStudy (anonymous):

\[120x - 2x^2\]

OpenStudy (anonymous):

Let's say the right side of the rectangle is the garage. The top width is length x. Therefore the bottom width is also length x. The left side will therefore be 120-2x in length. Area can be found through (120-2x)(x)=-2x(x+60)

OpenStudy (anonymous):

f(x) = (120-2x) (x)

OpenStudy (anonymous):

\[\text{area}=x \left(\frac{120-x}{2}\right)=\frac{1}{2} \left(120 x-x^2\right) \]